Find the price for good Z and the quantity supply for good X (show all the calculations) if: (i) The elasticity of supply is equal to 1 and the price increases from $40 to $50.

(ii) If the elasticity of demand is 0.5 and the quantity demanded decreases from 95.000 to 85.000.

(iii) Draw the graph and indicate the equilibrium price and quantity .


Price per Tonne ($) Quantity Demanded Quantity Supplied

40 150 80

50 120 X

60 110 110

80 95 115

Z 85 120

110 80 140


1
Expert's answer
2022-02-04T17:59:37-0500

(i)

"Es=\frac{\Delta Q}{\Delta P}"

Es=1


"\Delta P=\frac{50}{40}=1.25"


"1=\frac{\Delta Q}{1.25}"


"1.25=\frac{Q2}{80}"


Q2=100


(ii)


"0.5=\frac{0.89}{\Delta P}"


"\Delta P=1.78"


"1.78=\frac{P2}{80}"


P2=142.4


(iii)



P=60, Q=110